# Stein Variational Evolution Strategies

### v1

@@ 3.2 Stein Variational CMA-ES
Based on the evolutionary SVGD update in Eq. (10) and the CMA-ES update of the search distribution mean in Eq. 4, we now define Stein Variational CMA-ES (SV-CMA-ES). SV-CMA-ES is a multi-population version of CMA-ES, where $n$ search distributions are updated in parallel, each representing an SVGD particle $\mathbf{x}_{i}$ via their distribution mean. In other words, for each particle $\mathbf{x}_{i}$, there is a corresponding Gaussian search distribution that is centered at the particle and parametrized as ${\cal N}(\mathbf{x}_{i},\sigma_{i}^{2}\mathbf{C}_{i})$. This permits us to approximate the score function $\nabla_{\mathbf{x}}f(\mathbf{x})$ using the CMA-ES mean update in Eq. (4).
Given a particle $\mathbf{x}_{i}$ and a sampled subpopulation $\boldsymbol{\xi}_{ij}\sim{\cal N}(\mathbf{x}_{i},\sigma_{i}^{2}\mathbf{C}_{i})$, we approximate the SVGD update as follows:
$\mathbf{x}_{i}\leftarrow\mathbf{x}_{i}+\epsilon\operatorname*{\mathbb{E}}_{\mathbf{x}\sim q}\left[\tilde{\nabla}_{\mathbf{x}}f(\mathbf{x})k(\mathbf{x},\mathbf{x}_{i})+\nabla_{\mathbf{x}}k(\mathbf{x},\mathbf{x}_{i})\right]$
$\approx\mathbf{x}_{i}+\frac{\epsilon}{n}\sum_{j=1}^{n}\Biggl{[}\Bigl{[}\sigma_{j}\sum_{k=1}^{m}w_{jk}y_{jk}\Bigr{]}k(\mathbf{x}_{j},\mathbf{x}_{i})+\nabla_{\mathbf{x}_{j}}\underbrace{k(\mathbf{x}_{j},\mathbf{x}_{i})}_{k_{ji}}\Biggr{]}$
$=\mathbf{x}_{i}+\frac{\epsilon}{n}\sum_{j=1}^{n}\Biggl{[}\Bigl{[}\sum_{k=1}^{m}w_{jk}(\boldsymbol{\xi}_{jk}-\mathbf{x}_{j})\Bigr{]}k_{ji}+\nabla_{\mathbf{x}_{j}}k_{ji}\Biggr{]}$ (11)
where $w_{jk}$ are the sample weights that are computed based on the fitness values $f(\boldsymbol{\xi}_{i})$ following Hansen (2016). Note that in the last line, the SVGD learning rate hyperparameter $\epsilon$ can easily be replaced by the CMA-ES step-size hyperparameter $\alpha_{\mathbf{x}}$.
It now remains to define the step-size control for CMA-ES from which the covariance update naturally follows. The CMA-ES step-size update (7) is based on the length of the step $\hat{\mathbf{y}}$ by which the particle, i.e. distribution mean, is shifted. Given the particle update in Eq. (11), we define $\hat{\mathbf{y}}_{i}$:
$\hat{\mathbf{y}}_{i}=\sum_{j=1}^{n}\Biggl{[}\Bigl{[}\sum_{k=1}^{m}w_{jk}(\boldsymbol{\xi}_{jk}-\mathbf{x}_{j})\Bigr{]}k(\mathbf{x}_{j},\mathbf{x}_{i})+\nabla_{\mathbf{x}_{j}}k_{ji}\Biggr{]}$ (12)
@@ 3.3 Practical considerations
We now discuss some modifications to the algorithm that we found beneficial in practice. As noted earlier, the update of the particle in Eq. (11) smoothens the gradient approximation across all particles. As a result, the magnitude of the effective steps is reduced compared to standard CMA-ES. Since CMA-ES reduces the step-size $\sigma$ automatically when small steps are taken, this may lead to premature convergence. An example that illustrates this problem is a bimodal distribution with both modes far apart, such that for most particles $k(\mathbf{x},\mathbf{y})$ is close to zero for all pairs $\mathbf{x},\mathbf{y}$ that are sampled from different modes. In this scenario, the driving force term of the SVGD update corresponds to the vanilla CMA-ES update scaled down by the factor of $1/n$. Hence, the proposed steps in this scenario would shrink iteratively. To address this issue, we propose the following simplified particle update that we found to work well in practice:
@@
$\begin{split}&\mathbf{x}_{i}+\frac{\epsilon}{n}\sum_{j=1}^{n}\Biggl{[}\Bigl{[}\sum_{k=1}^{m}w_{ik}(\boldsymbol{\xi}_{ik}-\mathbf{x}_{i})\Bigr{]}+\nabla_{\mathbf{x}_{j}}k(\mathbf{x}_{j},\mathbf{x}_{i})\Biggr{]}.\\ \end{split}$ (13)
@@
This updates uses only particle $\mathbf{x}_{i}$ to estimate the first term of the update, i.e., the driving force. We note that this corresponds to a hybrid kernel SVGD (D’Angelo et al., 2021; MacDonald et al., 2023) step $\phi(\mathbf{x})=\operatorname*{\mathbb{E}}_{\mathbf{x}\sim q}\left[\nabla_{\mathbf{x}}f(\mathbf{x})k_{1}(\mathbf{x},\mathbf{x}_{i})+\nabla_{\mathbf{x}}k_{2}(\mathbf{x},\mathbf{x}_{i})\right]$, with $k_{1}(\mathbf{x},\mathbf{y})=n\mathds{1}(x=y)$. This kernel can be approximated by an RBF kernel with small bandwidth $h\rightarrow 0$. In the latter case, it follows from Theorem 4.1 in MacDonald et al. (2023) that Eq. (13) indeed approximates SVGD if $k_{2}$ is chosen to be an arbitrary RBF kernel.
@@
While the update in Eq. (13) does not possess the same capabilities of transporting particles “along a necklace” as the vanilla SVGD update (cf. Fig. 1 of Liu and Wang (2016)), it has been noted previously that these SVGD capabilities play a limited role for practical problems in the first place (D’Angelo and Fortuin, 2021). Instead, prior work proposed the annealed update in Eq. (3) to transport the particles to regions of high density (D’Angelo and Fortuin, 2021; Liu et al., 2017). In practice, we observe that by using an annealed version of the above update, i.e.
@@
$\mathbf{x}_{i}+\epsilon\Biggl{[}\sum_{k=1}^{m}w_{ik}(\boldsymbol{\xi}_{ik}-\mathbf{x}_{i})+\frac{\gamma(t)}{n}\sum_{j=1}^{n}\nabla_{\mathbf{x}_{j}}k(\mathbf{x}_{j},\mathbf{x}_{i})\Biggr{]}$ (14)
@@
encourages sufficient mode coverage to efficiently sample from distributions.
Algorithm 1 Stein Variational CMA-ES. Differences to the parallel CMA-ES algorithm are highlighted in red.
@@
Input: Kernel $k(\cdot,\cdot)$; elite subpop. size $\lambda$; elite weights $w_{i=1\dots m}$; learning rates $\alpha_{\mathbf{x}},\alpha_{\sigma},\alpha_{1},\alpha_{\lambda},\alpha_{c}$; damping hyperparameter $d_{\sigma}$; dimension $d$
@@
1:Initialize population parameters $\mathbf{x}_{i},\sigma_{i},\mathbf{C}_{i}$ for particles $i=1,\dots,n$
@@
$\boldsymbol{\xi}_{ij}\sim{\cal N}(\mathbf{x}_{i},\sigma_{i}^{2}\mathbf{C}_{i})\text{, for $j=1,\dots,m$}$
@@
5:     Estimate gradient & shift particle:
6:     Sort $\mathbf{y}_{ij}$ by $f_{ij}=f(\boldsymbol{\xi}_{ij})$ in ascending order
${\color[rgb]{0.9,0.17,0.31}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.17,0.31}\hat{\mathbf{y}}_{i}=\sum_{k=1}^{m}w_{ik}(\boldsymbol{\xi}_{ik}-\mathbf{x}_{i})+\frac{\gamma(t)}{n}\sum_{j=1}^{n}\nabla_{\mathbf{x}_{j}}k(\mathbf{x}_{j},\mathbf{x}_{i})}$
${\color[rgb]{0.9,0.17,0.31}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.17,0.31}\mathbf{x}_{i}\leftarrow\mathbf{x}_{i}+\alpha_{\mathbf{x}}\sigma_{i}~{}\hat{\mathbf{y}}_{i}}$
@@
$\lambda_{\text{eff}}=(\textstyle\sum_{\ell=1}^{\lambda}w_{i\ell}^{2})^{-1}$
$\begin{split}&\mathbf{p}_{\sigma_{i}}\leftarrow(1-\alpha_{\sigma})\mathbf{p}_{\sigma_{i}}\\ &\qquad+\sqrt{\alpha_{\sigma}(2-\alpha_{\sigma})~{}\lambda_{\text{eff}}}~{}\mathbf{C}_{i}^{-\frac{1}{2}}{\color[rgb]{0.9,0.17,0.31}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.17,0.31}\hat{\mathbf{y}}_{i}}\end{split}$
@@
$d(h_{\sigma})=1\text{ if }\alpha_{c}(1-h_{\sigma})(2-\alpha_{c})\leq 1\text{ else }0$
$\bar{w}_{ij}=w_{ij}\text{ if }w_{ij}\geq 0\text{ else }w_{ij}d/\lVert\mathbf{C}_{i}^{-\frac{1}{2}}\mathbf{y}_{ij}\rVert^{2}$
$\mathbf{p}_{i}^{c}\leftarrow(1-\alpha_{c})\mathbf{p}_{i}^{c}+h_{\sigma_{i}}\sqrt{\alpha_{c}(2-\alpha_{c})\lambda_{\text{eff}}}~{}{\color[rgb]{0.9,0.17,0.31}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.17,0.31}\hat{\mathbf{y}}_{i}}$
$\begin{split}&\mathbf{C}_{i}\leftarrow(1+d(h_{\sigma_{i}})-\alpha_{1}-\alpha_{\lambda}\textstyle\sum_{j=1}^{m}w_{ij})\mathbf{C}_{i}\\ &\qquad+\alpha_{1}\mathbf{p}_{i}^{c}{\mathbf{p}_{i}^{c}}^{T}+\alpha_{\lambda}\textstyle\sum_{j=1}^{m}\bar{w}_{ij}\mathbf{y}_{ij}\mathbf{y}_{ij}^{T}\end{split}$
@@ 4 Related Work
SVGD is a popular method for sampling from unnormalized densities. As such, SVGD has been an active research topic and many extensions have been proposed. These include approaches to improve the performance in high dimensions, for instance using projections (Chen et al., 2019) or by adjusting the particle update to reduce its bias (D’Angelo and Fortuin, 2021; Ba et al., 2021). Other extensions include non-Markovian steps (Ye et al., 2020; Liu et al., 2022), learning-based methods (Langosco di Langosco et al., 2021; Zhao et al., 2023), and domain-specific kernel functions (Sharma et al., 2023; Barcelos et al., 2024). While our focus lies on gradient-free SVGD approaches, most of these ideas could be integrated into our approach, which would be an interesting direction of future research.
@@
Many gradient-free sampling methods, like those in the MCMC family, iteratively update a proposal distribution to match the target (Andrieu et al., 2003). A shortcoming of these approaches is their slower sampling procedure compared to SVGD, as they often require many burn-in steps and additional rejection steps to decorrelate the samples. Gradient-free SVGD (Han and Liu, 2018, GF-SVGD) addresses this by estimating the gradient for SVGD on a surrogate distribution, which allows for between-chain interactions and requires no rejections. Further work improved the computational efficiency of this method by fitting the surrogate to a limited set of points (Yan and Zou, 2021). However, surrogate-based methods require a well-chosen prior to initialize the surrogate, as they lack an explicit exploration mechanism. Thus, in practical scenarios, a different gradient-free SVGD approach has been presented which relies on MC estimates of the gradient (Liu et al., 2017). In this work, we present a novel perspective on gradient-free SVGD, which combines ideas from the literature on ES and variational inference. Different to prior work, we propose a particle update that is based on CMA-ES, a highly efficient ES (Hansen and Kern, 2004).
@@
ES are a specific class of BBO methods that iteratively improve a search distribution over solution candidates by implementing specific sampling, evaluation and update mechanisms (Rechenberg, 1978). While ES commonly use a single distribution (Salimans et al., 2017; Wierstra et al., 2014; Sehnke et al., 2010), it has been demonstrated that their efficiency can be improved by employing restarts or multiple runs in parallel (Auger and Hansen, 2005; Pugh et al., 2016). For instance, restarts with increasing population sizes have been demonstrated to improve the performance of CMA-ES (Loshchilov et al., 2012). A downside of restarting approaches is their sequential nature, which makes them slower and prohibits to exploit the benefits of modern GPUs. Our method is different as it uses the SVGD update to sample multiple subpopulations in parallel which naturally enables to explore multiple modes. In particular, our proposed SVGD-based update is simpler to compute than other distributed updates (Wang et al., 2019b), yet more informed than uncoordinated parallel runs.

### v2

@@ 3 A Stein Variational Evolution Strategy
Based on the SVGD update in Eq. (2) and the CMA-ES update of the search distribution mean in Eq. (4), we now define Stein Variational CMA-ES (SV-CMA-ES). The full algorithm is listed in Algorithm 1. SV-CMA-ES is a multi-population version of CMA-ES, where $\varrho$ search distributions are updated in parallel, each representing an SVGD particle $\mathbf{x}_{i}$ via their distribution mean. In other words, for each particle, there is a corresponding Gaussian search distribution that is centered at the particle and parametrized as ${\cal N}(\mathbf{x}_{i},\sigma_{i}^{2}\mathbf{C}_{i})$. Given the standard CMA-ES distribution update step $\Delta\mathbf{x}_{\textsc{cma}}$ from Eq. (5) and a sampled population $\boldsymbol{\xi}_{ij}\sim{\cal N}(\mathbf{x}_{i},\sigma_{i}^{2}\mathbf{C}_{i})$, we propose the following SVGD-based update:
$\mathbf{x}_{i}$ $\leftarrow\mathbf{x}_{i}+\epsilon\leavevmode\nobreak\ \phi(\mathbf{x}_{i})\quad\text{with}$ (11)
$\phi(\mathbf{x}_{i})$ $=\operatorname*{\mathbb{E}}_{\mathbf{x}_{j}\sim q}\Bigl{[}k(\mathbf{x}_{j},\mathbf{x}_{i})\Delta{\mathbf{x}_{j}}_{\textsc{cma}}+\nabla_{\mathbf{x}_{j}}k(\mathbf{x}_{j},\mathbf{x}_{i})\Bigr{]}$
@@
where we assume the same sorting by fitness in our sum as in vanilla CMA-ES and $w_{j\ell}$ are the sample weights that are computed based on the fitness values $f(\boldsymbol{\xi}_{j\ell})$ following Hansen (2016). Further, we use an additional step-size hyperparameter $\epsilon$ for notational consistency with SVGD, but we always fix it to $\epsilon=1$.
Eq. (12) defines how to update each particle search distribution mean. It now remains to define the remaining SV-CMA-ES parameter updates. The original CMA-ES step-size update (8) is based on the length of the distribution mean update step. In the particle update in Eq. (11), this quantity corresponds to the effective update step $\phi(\mathbf{x}_{i})$. Given this particle shift, the smoothened step estimate $\mathbf{p}_{\sigma_{i}}$ is computed analogously to the CMA-ES optimization path update in Eq. (7):
$\mathbf{p}_{\sigma_{i}}\leftarrow(1-\alpha_{\sigma})\mathbf{p}_{\sigma_{i}}+\sqrt{\alpha_{\sigma}(2-\alpha_{\sigma})\leavevmode\nobreak\ m_{\text{eff},i}}\leavevmode\nobreak\ \mathbf{C}_{i}^{-\frac{1}{2}}\phi(\mathbf{x}_{i})/\sigma_{i}$ (13)
Using the same construction, we update $\mathbf{p}_{c_{i}}$ based on $\phi(\mathbf{x}_{i})$, from which the covariance $\mathbf{C}_{i}$ can be computed using Equation 9.
@@ 3.1 Practical considerations
We now discuss some modifications to the algorithm that we found beneficial in practice. As noted earlier, the update of the particle in Eq. (12) smoothens the gradient approximation across all particles. As a result, the magnitude of the effective steps is reduced compared to standard CMA-ES. Since CMA-ES reduces the step-size $\sigma$ automatically when small steps are taken, this may lead to premature convergence. An example that illustrates this problem is a bimodal distribution with both modes far apart, such that for most particles $k(\mathbf{x},\mathbf{y})$ is close to zero for all pairs $\mathbf{x},\mathbf{y}$ that are sampled from different modes. In this scenario, the driving force term of the update corresponds to the vanilla CMA-ES update, scaled down by the factor of $1/\varrho$. Hence, the proposed steps in this scenario would shrink iteratively. To address this issue, we propose the following simplified particle update:
@@
$\begin{split}&\phi(\mathbf{x}_{i})=\frac{1}{\varrho}\sum_{j=1}^{\varrho}\Biggl{[}\Bigl{[}\sum_{\ell=1}^{m}w_{i\ell}(\boldsymbol{\xi}_{i\ell}-\mathbf{x}_{i})\Bigr{]}+\nabla_{\mathbf{x}_{j}}k(\mathbf{x}_{j},\mathbf{x}_{i})\Biggr{]}.\\ \end{split}$ (14)
@@
This update uses only the particle $\mathbf{x}_{i}$ to estimate the first term of the update, i.e., the driving force. We note that this corresponds to a hybrid kernel SVGD setting (D’Angelo et al., 2021; MacDonald et al., 2023), which uses two separate kernels to compute the repulsion and driving force terms: $\phi_{\text{hybrid}}(\mathbf{x}_{i})=\operatorname*{\mathbb{E}}_{\mathbf{x}\sim q}\left[\nabla_{\mathbf{x}}f(\mathbf{x})k_{1}(\mathbf{x},\mathbf{x}_{i})+\nabla_{\mathbf{x}}k_{2}(\mathbf{x},\mathbf{x}_{i})\right]$ if we choose $k_{1}(\mathbf{x},\mathbf{y})=n\mathds{1}(\mathbf{x}=\mathbf{y})$. This kernel can be approximated by an RBF kernel with small bandwidth $h\rightarrow 0$.
@@
While the update in Eq. (14) does not possess the same capabilities of transporting particles “along a necklace” as the vanilla SVGD update (cf. Fig. 1 of Liu and Wang (2016)), it has been noted that these SVGD capabilities play a limited role for practical problems in the first place (D’Angelo and Fortuin, 2021). Instead, prior work proposed the annealed update in Eq. (3) to transport the particles to regions of high density (D’Angelo and Fortuin, 2021; Liu et al., 2017). In practice, we observe that using the annealed version of the above update, i.e.,
@@
$\phi(\mathbf{x}_{i})$ $=\frac{1}{\varrho}\sum_{j=1}^{\varrho}\Biggl{[}\Bigl{[}\sum_{\ell=1}^{m}w_{i\ell}(\boldsymbol{\xi}_{i\ell}-\mathbf{x}_{i})\Bigr{]}+\gamma(t)\nabla_{\mathbf{x}_{j}}k(\mathbf{x}_{j},\mathbf{x}_{i})\Biggr{]}$
$=\sum_{\ell=1}^{m}w_{i\ell}(\boldsymbol{\xi}_{i\ell}-\mathbf{x}_{i})+\frac{\gamma(t)}{\varrho}\sum_{j=1}^{\varrho}\nabla_{\mathbf{x}_{j}}k(\mathbf{x}_{j},\mathbf{x}_{i})$ (15)
@@
ensures sufficient mode coverage to efficiently sample from distributions.
The substitution of the score function with the CMA-ES step introduces a bias in comparison to the SVGD update in Eq. (3), meaning it does not strictly adhere to the canonical SVGD framework and does not inherit its robust convergence properties. Still, we find that, in practice, the update makes a useful tradeoff which combines the computational efficiency of CMA-ES with the particle set entropy preservation capabilities of SVGD. We leave a more in-depth theoretical analysis of the algorithm for future work, and present our empirical findings in the subsequent section. For an empirical convergence analysis, we refer to Appendix C.4.
Algorithm 1 Stein Variational CMA-ES. Differences to the parallel CMA-ES algorithm are highlighted in blue.
@@
Input: Kernel $k(\cdot,\cdot)$; num. particles $\varrho$; subpop. size $n$; num. elites $m$; elite weights $w_{i=1\dots m}$; learning rates $\epsilon,\alpha_{\sigma},\alpha_{1},\alpha_{m},\alpha_{c}$; damping hyperparam. $d_{\sigma}$; dimension $d$; num. iterations $T$
@@
1:Initialize population parameters $\mathbf{x}_{i},\sigma_{i},\mathbf{C}_{i}$ for particles $i=1,\dots,\varrho$
@@
$\boldsymbol{\xi}_{ij}\sim{\cal N}(\mathbf{x}_{i},\sigma_{i}^{2}\mathbf{C}_{i})\text{, for $j=1,\dots,n$}$
@@
5:     Estimate gradient & shift particle (Eq. 15):
6:     Sort samples by $f_{ij}=f(\boldsymbol{\xi}_{ij})$ in ascending order
${\color[rgb]{0.0703125,0.36328125,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.36328125,0.65234375}\phi(\mathbf{x}_{i})=\sum_{\ell=1}^{m}w_{i\ell}(\boldsymbol{\xi}_{i\ell}-\mathbf{x}_{i})+\frac{\gamma(t)}{\varrho}\sum_{j=1}^{\varrho}\nabla_{\mathbf{x}_{j}}k(\mathbf{x}_{j},\mathbf{x}_{i})}$
${\color[rgb]{0.0703125,0.36328125,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.36328125,0.65234375}\mathbf{x}_{i}\leftarrow\mathbf{x}_{i}+\epsilon\phi(\mathbf{x}_{i})}$
@@
$m_{\text{eff},i}=(\textstyle\sum_{\ell=1}^{m}w_{i\ell}^{2})^{-1}$
$\begin{split}&\mathbf{p}_{\sigma_{i}}\leftarrow(1-\alpha_{\sigma})\mathbf{p}_{\sigma_{i}}\\ &\qquad+\sqrt{\alpha_{\sigma}(2-\alpha_{\sigma})\leavevmode\nobreak\ m_{\text{eff},i}}\leavevmode\nobreak\ \mathbf{C}_{i}^{-\frac{1}{2}}{\color[rgb]{0.0703125,0.36328125,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.36328125,0.65234375}\phi(\mathbf{x}_{i})/\sigma_{i}}\end{split}$
@@
$d(h_{\sigma_{i}})=1\text{ if }\alpha_{c}(1-h_{\sigma})(2-\alpha_{c})\leq 1\text{ else }0$
$\bar{w}_{ij}=w_{ij}\text{ if }w_{ij}\geq 0\text{ else }d/\lVert\mathbf{C}_{i}^{-\frac{1}{2}}\mathbf{y}_{ij}\rVert^{2}$
$\mathbf{p}_{c_{i}}\!\leftarrow\!(1-\alpha_{c})\mathbf{p}_{c_{i}}\!+\!h_{\sigma_{i}}\sqrt{\alpha_{c}(2\!-\!\alpha_{c})m_{\text{eff},i}}\leavevmode\nobreak\ {\color[rgb]{0.0703125,0.36328125,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.36328125,0.65234375}\phi(\mathbf{x}_{i})/\sigma_{i}}$
$\begin{split}&\mathbf{C}_{i}\leftarrow(1+d(h_{\sigma_{i}})-\alpha_{1}-\alpha_{m}\textstyle\sum_{j=1}^{n}w_{ij})\mathbf{C}_{i}\\ &\qquad+\alpha_{1}\mathbf{p}_{c_{i}}\mathbf{p}_{c_{i}}^{T}+\alpha_{m}\textstyle\sum_{j=1}^{n}\bar{w}_{ij}\mathbf{y}_{ij}\mathbf{y}_{ij}^{T}\end{split}$
@@ 4 Related Work
SVGD is a popular method for sampling from unnormalized densities. As such, SVGD has been an active field of research and many extensions have been proposed. These include approaches to improve the performance in high dimensions, for instance using projections (Chen et al., 2019) or by adjusting the particle update to reduce its bias (D’Angelo and Fortuin, 2021; Ba et al., 2021). Other extensions include non-Markovian steps (Ye et al., 2020; Liu et al., 2022), learning-based methods (Langosco et al., 2021; Zhao et al., 2023), and domain-specific kernel functions (Sharma et al., 2023; Barcelos et al., 2024). While our focus lies on gradient-free SVGD approaches, most of these ideas could be integrated into our approach, which would be an interesting direction of future research.
@@
Many gradient-free sampling methods, like those in the MCMC family, iteratively update a proposal distribution to match the target (Andrieu et al., 2003). A shortcoming of these approaches is their slower sampling procedure compared to SVGD, as they are prone to be trapped in a single mode over long periods of time on multimodal objectives. Population-based MCMC methods improve over this by running multiple chains in parallel, which exchange information over time (Laskey and Myers, 2003). Notably, parallel tempering methods simulate chains with different temperatures in parallel to improve mode coverage (Swendsen and Wang, 1986). Still, these methods commonly require sample rejections and potentially long burning-in periods. Gradient-free SVGD (Han and Liu, 2018, GF-SVGD) addresses this by estimating the gradient for SVGD on a surrogate distribution, which allows for interactions between all chains at each update step and fast convergence rates. Further work improved the computational efficiency of this method by fitting the surrogate to a limited set of points (Yan and Zou, 2021). However, these surrogate-based methods require a well-chosen prior for surrogate initialization, as they lack an explicit exploration mechanism. Thus, in practical scenarios, a different gradient-free SVGD approach has been presented which relies on MC estimates of the gradient (Liu et al., 2017). In this work, we present a novel perspective on gradient-free SVGD, which combines ideas from the literature on ES. Different from prior work, we propose a particle update that is based on CMA-ES, a highly efficient ES (Hansen and Kern, 2004).
@@
ES are a specific class of blackbox optimization methods that iteratively improve a search distribution over solution candidates by implementing specific sampling, evaluation and update mechanisms (Rechenberg, 1978). While ES commonly use a single distribution (Li et al., 2020), it has been demonstrated that their efficiency can be improved by employing restarts or multiple runs in parallel (Auger and Hansen, 2005; Pugh et al., 2016). For instance, restarts with increasing population sizes have been demonstrated to improve the performance of CMA-ES (Loshchilov et al., 2012). A downside of restarting approaches is their sequential nature, which makes them slower and prohibits exploiting the benefits of modern GPUs. Our method is different as it uses the SVGD update to sample multiple subpopulations in parallel, which naturally enables to explore multiple modes. In particular, our proposed SVGD-based update is simpler to compute than other distributed updates (Wang et al., 2019b), yet more informed than uncoordinated parallel runs.